OpenStudy (pulsified333):

For this problem, assume the balls in the box are numbered 1 through 8, and that an experiment consists of randomly selecting 3 balls one after another without replacement. There are 336 possible outcomes. What probability should be assigned to the event that at least one ball has an odd number?

OpenStudy (anonymous):

Would it be approximately .93

OpenStudy (pulsified333):

@emily_wilson how did you get that? It has to be as a fraction. I'm sorry I didn't specify

OpenStudy (anonymous):

52/56=13/14=.93

OpenStudy (pulsified333):

Thank you very much

OpenStudy (anonymous):

So it helped? You're welcome & Thank you!

OpenStudy (pulsified333):

Yes it did

OpenStudy (anonymous):

Glad I could help!!